# Imaging nuclei by smashing them at high energies: how are their shapes revealed after destruction?

### v2

@@ Abstract
Imaging-by-smashing also offers new experimental insights into subbarrier nuclear fusion reactions, a quantum tunneling phenomenon critically influenced by the Coulomb barrier Hagino and Takigawa (2012). This barrier is sensitive to the orientation of colliding nuclei and the interplay between reaction timescale and timescales associated with nuclear deformations. Specifically, static nuclear deformations, characterized by longer rotational timescales, directly affect fusion probability, whereas dynamic shape fluctuations arising from nuclear vibrations, with their inherently shorter timescales, may have a smaller impact on the Coulomb barrier. The unique capability of high-energy collisions to differentiate between static deformation and dynamic shape fluctuations is therefore vital. This distinction may help deciphering why sub-barrier fusion cross sections are significantly modified and to determining whether static or dynamic nuclear properties exert a more important influence on this quantum tunneling process.
@@
This method can also help address fundamental questions in nuclear physics. For instance, nuclear shape could generate nuclear electric dipole moment (EDM) though the laboratory Shiff moment that is proportional to the product of the quadrupole deformation parameter and square of the octupole deformation parameter $\beta_{3}^{2}\beta_{2}$ Spevak et al. (1997). One could use specially designed observables, such as correlation between trangular flow and radial flow to constrain the Shiff moment, $\left\langle v_{3}^{2}\delta[p_{\mathrm{T}}]\right\rangle\approx a+b\beta_{3}^{2}\beta_{2}$ ($\beta_{3}^{2}$ arises from $v_{3}^{2}$, while $\beta_{2}$ from $\delta[p_{\mathrm{T}}]$). 238U, though not a direct candidate for atomic EDM searches due to its even-even nature, could be used to characterize nuclear properties like its large static $\beta_{2}$ and a signficant $\beta_{3}$ fluctuation STAR Collaboration (2025) that are relevant to the Schiff moment, thus informing EDM measurements in odd-mass nuclei. The method can also used to reduce theoretical uncertainties in neutrinoless double-beta decay searches ($0\nu\beta\beta$) Agostini et al. (2023), a process highly sensitive to the similarity of nuclear shapes in parent-daughter isobaric pairs. By comparing collisions of these nuclei, the technique can precisely quantify shape differences, thereby improving decay rate predictions. Furthermore, it opens new avenues for studying alpha clustering in light isobar-like systems such as 16O+16O and 20Ne+20Ne Giacalone et al. (2024), providing stringent tests for ab initio nuclear theories.
@@
Conclusion. The imaging-by-smashing technique represents a paradigm shift in nuclear structure studies, transforming the apparent destructiveness of high-energy collisions into a powerful imaging tool. By exploiting the short timescales and collective dynamics of QGP formation and evolution, this method accesses nuclear many-body wavefunctions in ways inaccessible through conventional techniques. This method offers exciting possibilities for exploring shape coexistence, nuclear fusion, testing ab initio nuclear theories, and connecting nuclear structure to fundamental symmetries. The convergence of ultra-relativistic heavy-ion collisions with precision nuclear structure physics opens new frontiers that neither field could access independently. As we refine calibration procedures and expand to new collision systems, imaging-by-smashing promises to illuminate fundamental aspects of nuclear structure while advancing our understanding of QGP properties.
Acknowledgments. The author would like to thank Aihong Tang for inviting him to present a BNL colloquium, which forms the basis of this note, as well as critial feedback from Giuliano Giacalone. This work is supported by the U.S. Department of Energy, Award number DE-SC0024602.

### v4

@@ Abstract
Imaging-by-smashing also offers new experimental insights into subbarrier nuclear fusion reactions, a quantum tunneling phenomenon critically influenced by the Coulomb barrier Hagino and Takigawa (2012). This barrier is sensitive to the orientation of colliding nuclei and the interplay between reaction timescale and timescales associated with nuclear deformations. Specifically, static nuclear deformations, characterized by longer rotational timescales, directly affect fusion probability, whereas dynamic shape fluctuations arising from nuclear vibrations, with their inherently shorter timescales, may have a smaller impact on the Coulomb barrier. The unique capability of high-energy collisions to differentiate between static deformation and dynamic shape fluctuations is therefore vital. This distinction may help decipher why sub-barrier fusion cross sections are significantly modified and determine whether static or dynamic nuclear properties exert a more important influence on this quantum tunneling process.
@@
This method may also help address fundamental questions in nuclear physics. For instance, nuclear shape could generate nuclear electric dipole moment (EDM) through the laboratory Schiff moment that is proportional to the product of the quadrupole deformation parameter and square of the octupole deformation parameter $\beta_{3}^{2}\beta_{2}$ Spevak et al. (1997). One may use specially designed observables, such as correlation between triangular flow and radial flow, to constrain the Schiff moment, $\left\langle v_{3}^{2}\delta[p_{\mathrm{T}}]\right\rangle\approx a-b\beta_{3}^{2}\beta_{2}$ (where $\beta_{3}^{2}$ arises from $v_{3}^{2}$, while $\beta_{2}$ arises from $\delta[p_{\mathrm{T}}]$). 238U, though not a direct candidate for atomic EDM searches due to its even-even nature, could be used to characterize nuclear properties like its large static $\beta_{2}$ and a significant $\beta_{3}$ fluctuation STAR Collaboration (2025) that are relevant to the Schiff moment, thus informing EDM measurements in odd-mass nuclei. The method may also be used to reduce theoretical uncertainties in neutrinoless double-beta decay searches ($0\nu\beta\beta$) Agostini et al. (2023), a process highly sensitive to the similarity of nuclear shapes in parent-daughter isobaric pairs. By comparing collisions of these nuclei, the technique can precisely quantify shape differences, thereby improving decay rate predictions. Furthermore, it opens new avenues for studying alpha clustering in light isobar-like systems such as 16O+16O and 20Ne+20Ne Giacalone et al. (2025), providing stringent tests for ab initio nuclear theories.
@@
Conclusion. The imaging-by-smashing technique demonstrates the potential to turn the apparent destructiveness of high-energy collisions into a complementary imaging approach for nuclear structure studies. By exploiting the short timescales and collective dynamics of QGP formation and evolution, this method accesses nuclear many-body wavefunctions in ways that complement insights from conventional techniques. However, the successful application of this method requires major theoretical and experimental developments. Nevertheless, once properly validated, imaging-by-smashing could offer exciting possibilities, such as exploring shape coexistence, nuclear fusion, testing ab initio nuclear theories, and connecting nuclear structure to fundamental symmetries. I look forward to the successful development and validation of these techniques, and the eventual convergence of ultra-relativistic heavy-ion collisions with nuclear structure physics.
Acknowledgments. The author would like to thank Aihong Tang for inviting him to present a BNL colloquium, which forms the basis of this note, as well as for valuable discussions with Peter Butler, Giuliano Giacalone, Kouichi Hagino, and John Woods. This work is supported by the U.S. Department of Energy, Award number DE-SC0024602.